What volume, in mL, of 1.68 M H2SO4(aq) is needed to COMPLETELY NEUTRALIZE 170. mL of 1.07 M NaOH(aq)?

50.5 mL

1.08×102 mL

54.1 mL

22.8 mL

2.17×102 mL
Answer #1

Balanced chemical equation is:
2 NaOH + H2SO4 ---> Na2SO4 + H2O


Here:
M(NaOH)=1.07 M
M(H2SO4)=1.68 M
V(NaOH)=170.0 mL

According to balanced reaction:
1*number of mol of NaOH =2*number of mol of H2SO4
1*M(NaOH)*V(NaOH) =2*M(H2SO4)*V(H2SO4)
1*1.07 M *170.0 mL = 2*1.68M *V(H2SO4)
V(H2SO4) = 54.1 mL

Answer: 54.1 mL